OpenStudy (anonymous):

I would like to create a rectangular orchid garden that abuts my house so that the house itself forms the northern boundary. The fencing for the southern boundary costs $20 per foot, and the fencing for the east and west sides costs $10 per foot. If I have a budget of $160 for the project, what are the dimensions of the garden with the largest area I can enclose? ] __ ft(smaller value) ___ ft (larger value) ___ft^2

OpenStudy (radar):

Let x = length of southern boundary let y= length of east/west boundary. Now calculate costs which is to sum up to $160 20x+20y = $160 divide thru by 20 getting x + y = 8 x=8 - y Now the area will be x*y A=(8-y)(y) A=8y-y^2 now compute A' , A'=8-2y for max let A'=0 0=8-2y 2y=8 y=4 x=8-y=4 Check out cost 4($20) + 8($10)= $160 Those are dimensions 32 sq ft. Dont' know what the minimum would be that seems like a silly question. Maybe somone will compute the minimum area and spend $160.

OpenStudy (radar):

Error alert. Area will be width x length or 4 X 4 = 16 sq ft.

OpenStudy (radar):

Not 32 sq ft.